cullenn:

heyy i need help w math lol, also theres a few problems i js have to post them one after each other so like i cant show the next question until this one is answered-

cullenn:

1 attachment
Puffer:

I suck at math my best advice is use ai

cullenn:

@puffer wrote:
I suck at math my best advice is use ai
then why respond;n;

adreonnaRaeee:

36

cullenn:

@adreonnaraeee wrote:
36
well i finished it but couldnt u have explained that-

Luigi0210:

In the beginning of the question you are given that \(\Delta KLM\) ~ \( \Delta KNO \) So you can set it up as a proportion \(\large \frac{KM}{KO} = \frac{KL}{KN}\)

cullenn:

@luigi0210 wrote:
In the beginning of the question you are given that \(\Delta KLM\) ~ \( \Delta KNO \) So you can set it up as a proportion \(\large \frac{KM}{KO} = \frac{KL}{KN}\)
tysmm

gelphielvr:

@puffer wrote:
I suck at math my best advice is use ai
Please don't advise people to use ai for math problems. Ai is often inaccurate in math, especially when It involves drawn figures. It wont help anyone if the person who posted this gets the wrong answer from ai. If you don't know the answer to a problem either ignore it or tag someone who you know will be able to help

Puffer:

Ok mb guys was just trying to be helpful

OnenOnlyy:

@cullenn wrote:
@adreonnaraeee wrote:
36
well i finished it but couldnt u have explained that-
u needa explain it nxt time